# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 5.2 SQ Lower Bounds for Learning Multiclass Linear Classifiers with RCN
Assume that there exists an SQ algorithm $\mathcal{A}$ that if given access to a distribution $D^{\prime}$ over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is standard normal returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[h(\mathbf{x})\neq y]\leq\mathrm{OPT}+\gamma/4$. Using $\mathcal{A}$, we can reduce the testing problem to the learning problem.
Note that by running $\mathcal{A}$ to distribution $D_{\mathbf{P}}\in\mathcal{D}$, $\mathcal{A}$ returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{\mathbf{P}}}[h(\mathbf{x})\neq y]\leq\mathrm{OPT}+\gamma/4$. However, running $\mathcal{A}$ with SQ oracle access to $D$ would result in a hypothesis $h$ that cannot perform better than $1-1/K,$ as $D$ is a product distribution with the uniform distribution over the labels. Since $\mathrm{OPT}<1-1/K-\gamma/2$, because $\mathbf{H}_{i,i}\geq 1/K+\gamma/2$ as $\sum_{j}\mathbf{H}_{i,j}=1,i\in[K]$, we have that with an additional query of accuracy less than $\gamma/4$, we can solve $\mathcal{B}(\mathcal{D},D)$. ∎
For $K$ not divisible by $4$, we add $r=4\mod K$ additional classes whose weight vectors are identical to $\mathbf{w}^{(1)}$, and with no noise, i.e., for these added classes $j$, we have that $\mathbf{H}_{jj}=1$ and $\mathbf{H}_{j,i}=\mathbf{H}_{i,j}=0$ for all $j\neq i$. For the remaining classes, i.e., for $i,j\in[K-r]$, we consider a matrix $\mathbf{H}_{i,j}$ that satisfies the conditions of Lemma 5.4. Note that since these additional classes have zero probability (as ties are broken by the smallest index), our SQ hardness result extends to the general case of any $K$. ∎
@@ Appendix B Omitted Content from Sections 2, 3 and 4
Appendix C Basics on Statistical Query Model
@@
Let $0<a,c<\frac{1}{2}$ and $m,n\in\mathbb{Z}_{+}$ such that $m\leq n^{a}$. There exists a set $S$ of $\Omega(n^{c})$ matrices in $\mathbb{R}^{m\times n}$ such that every $\mathbf{U}\in S$ satisfies $\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}_{m}$ and every pair $\mathbf{U},\mathbf{V}\in S$ with $\mathbf{U}\neq\mathbf{V}$ satisfies $\|\mathbf{UV}^{\top}\|_{F}\leq O(n^{2c-1+2a})$.

### v2

@@ 5.2 SQ Lower Bounds for Learning Multiclass Linear Classifiers with RCN
Assume that there exists an SQ algorithm $\mathcal{A}$ that if given access to a distribution $D^{\prime}$ over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is the standard normal returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[h(\mathbf{x})\neq y]\leq\mathrm{OPT}+\gamma/4$. Using $\mathcal{A}$, we can solve the testing problem.
Note that by running $\mathcal{A}$ to distribution $D_{\mathbf{P}}\in\mathcal{D}$, $\mathcal{A}$ returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{\mathbf{P}}}[h(\mathbf{x})\neq y]\leq\mathrm{OPT}+\gamma/4$. Moreover, by running $\mathcal{A}$ with SQ oracle access to $D$ results in a hypothesis $h$ that cannot perform better than $1-1/K,$ as $D$ is a product distribution with the uniform distribution over the labels. Since $\mathrm{OPT}<1-1/K-\gamma/2$, because $\mathbf{H}_{i,i}\geq 1/K+\gamma/2$ as $\sum_{j}\mathbf{H}_{i,j}=1,i\in[K]$, we have that with an additional query of tolerance less than $\gamma/4$, we can solve $\mathcal{B}(\mathcal{D},D)$. ∎
For $K$ not divisible by $4$, we add $r=4\mod K$ additional classes whose weight vectors are identical to $\mathbf{w}^{(1)}$, and with no noise, i.e., for these added classes $j$, we have that $\mathbf{H}_{j,j}=1$ and $\mathbf{H}_{j,i}=\mathbf{H}_{i,j}=0$ for all $j\neq i$. For the remaining classes, i.e., for $i,j\in[K-r]$, we consider a matrix $\mathbf{H}_{i,j}$ that satisfies the conditions of Lemma 5.5. Note that since these additional classes have zero probability (as ties are broken by the smallest index), our SQ hardness result extends to the general case of any $K$. This completes the proof. ∎
@@ Appendix A Additional Preliminaries
Legendre Polynomials
Legendre polynomials are orthogonal on the interval $[-1,1]$. Their key properties are summarized in the following fact.
Fact A.8 ([Sze67]).
The Legendre polynomials $P_{k}$ for $k\in\mathbb{Z}$, satisfy the following properties:
  1. 1.
    $P_{k}$ is a $k$-degree polynomial and $P_{0}(x)=1$ and $P_{1}(x)=x$.
  2. 2.
    $\int_{-1}^{1}P_{i}(x)P_{j}(x)\mathrm{d}x=2/(2i+1)\mathds{1}(i=j)$, for all $i,j\in\mathbb{Z}$.
  3. 3.
    $|P_{k}(x)|\leq 1$ for all $|x|\leq 1$.
  4. 4.
    $P_{k}(x)=(-1)^{k}P_{k}(-x)$.
  5. 5.
    $P_{k}(x)=2^{-k}\sum_{i=1}^{\lceil k/2\rceil}\binom{k}{i}\binom{2k-2i}{k}x^{k-2i}$.
@@ Appendix B Omitted Content from Sections 2, 3 and 4
Appendix C Additional Background on the SQ Model
@@
Let $0<a,c<\frac{1}{2}$ and $m,n\in\mathbb{Z}_{+}$ such that $m\leq n^{a}$. There exists a set $S$ of $2^{\Omega(n^{c})}$ matrices in $\mathbb{R}^{m\times n}$ such that every $\mathbf{U}\in S$ satisfies $\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}_{m}$ and every pair $\mathbf{U},\mathbf{V}\in S$ with $\mathbf{U}\neq\mathbf{V}$ satisfies $\|\mathbf{UV}^{\top}\|_{F}\leq O(n^{2c-1+2a})$.